OpenStudy (anonymous):

The sum of 3 consecutive ODD integers is 189. What is the first integer?

OpenStudy (anonymous):

Consecutive odd integers are x, x+2, x+ 4 so what equation would you write?

OpenStudy (anonymous):

well??

OpenStudy (anonymous):

Hint: SUM means to add up

OpenStudy (anonymous):

(2x + 1) + (2x + 3) + (2x + 5) = 189

OpenStudy (anonymous):

@Easyaspi314

OpenStudy (anonymous):

ok...now solve for x.

OpenStudy (anonymous):

X=30

OpenStudy (anonymous):

@Easyaspi314

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

so it will be 30

OpenStudy (anonymous):

x = 30. So the first integer will be 2x + 1 = 2(30) +1 = 61 61 is the first integer

OpenStudy (anonymous):

The question was to find the first integer. Your first integer is 2x+1...since you solved for x = 30, substitute into 2x+1 to get 61.

OpenStudy (anonymous):

oh okay, makes more since now, thank you for your help

OpenStudy (anonymous):

Welcome.

OpenStudy (anonymous):

63

OpenStudy (anonymous):

61 is the first integer.

OpenStudy (anonymous):

but 63+63+63=189?

OpenStudy (anonymous):

listen...you want 3 consecutive odd integers..63,63,63 are NOT 3 consecutive odd integers 61,63,65 are.

OpenStudy (anonymous):

oh i see. i misread the consecutive

OpenStudy (anonymous):

i wish i could listen